複素力学系と諸関連分野の総合的研究

研 究 集 会

京都大学数理解析研究所の共同事業のひとつとして，下記のように研究集会を催しますので，ご案内申し上げます．

研究代表者 木坂 正史

京都大学大学院人間環境学研究科研究副代表者 諸澤 俊介高知大学理学部数学科

記

日時：１９９９年１０月５日 (火) １０：００− １０月８日 (金) １２：１５場所：京都大学数理解析研究所４階４２０号室京都市左京区北白川追分町市バス 農学部前 または 北白川 下車

プログラム

１０月５日（火）１０：００ー１１：００ 前 川 和 俊 （京大 人間・環境） Shift-like mapping of dynamical degree golden ratio of C^{3} １１：１５ー１２：１５ 宇 敷 重 広 （京大 人間・環境） Fredholm determinant of complex Ruelle operator, zeta function and forward/backward Collet-Eckmann condition １３：３０ー１４：３０ 宍 倉 光 広 （広島大 理）射影空間上の正則力学系について１４：４５ー１５：４５ 糸 健 太 郎 （東工大 理） Action of mapping class group on extended Bers slice １６：００ー１７：００ Yair N. Minsky （State University of New York） Classification and rigidity problems for Kleinian groups

１０月６日（水）１０：００ー１１：００ 宮 地 秀 樹 （阪市大 理） On cusps in the boundary of the Maskit slice for once punctured torus groups １１：１５ー１２：１５ Yair N. Minsky （State University of New York） Geometry of punctured-torus groups １３：３０ー１４：３０ Yair N. Minsky （State University of New York） Curves, continued fractions and geometry, in the punctured torus and elsewhere １４：４５ー１５：４５ 小 森 洋 平 （阪市大 理） On holomorphic slices of the space of punctured torus groups １６：００ー１７：００ 須 川 敏 幸 （京大 理） 小 森 洋 平 （阪市大 理） Visualizing the Bers embedding of the Teichmüller space of a once-punctured torus

１０月７日（木）１０：００ー１１：００ 稲 生 啓 行 （京大 理） Rigidity of infinitely renormalizable polynomials of higher degree １１：１５ー１２：１５ 沈 維 考 （東大 数理） On the qc rigidity of real polynomials １３：３０ー１４：３０ 神 貞 介 （東大 数理） Dynamics of polynomial automorphisms of C^{2} １４：４５ー１５：４５ 中 根 静 男 （東京工芸大） Stretching rays for the family of real cubic polynomials １６：００ー１７：００ 角 大 輝 （東工大 理） Fiberwise Julia sets of rational skew products

１０月８日（金）１０：００−１１：００ 諸 澤 俊 介 （高知大 理） Appearance of Baker domains and wandering domains １１：１５ー１２：１５ 谷 口 雅 彦 （京大 理）構造的に有限な整函数

以 上